What is the rate of return for an investor who purchased $1,054.47 for a five-year bond with a 7% coupon and sells the bond one year later for $1,040.67?

A) 5.00%

B) 5.33%

C) 6.40%

D) 7.00%
Answer #1

B) 5.33%

Rate of return = (C+(P1-P0))/P0 Where,
= (70+(1040.67-1054.47))/1054.47 C = $         70.00
= 5.33% P0 = $ -1,054.47
P1 = $   1,040.67